OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x(x-1)=3 Step by step please, thanks!

TheSmartOne (thesmartone):

First multiply 2x(x-1) out ad then make it equal to 0 Finally use the quadratic formula and solve for it.

OpenStudy (jdoe0001):

if you expand 2x(x-1) = 3 what would it give you?

OpenStudy (anonymous):

2x^2-2x=3 2x^2-2x-3=0 Is that the right equation?

TheSmartOne (thesmartone):

Yes...

OpenStudy (jdoe0001):

yes \(\bf \textit{quadratic formula}\\ {\color{blue}{ 2}}x^2{\color{red}{ -2}}x{\color{green}{ -3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

Wow, I can't believe I didn't get that! It must have been one of those problems when you just go blank. Thanks guys!

OpenStudy (jdoe0001):

yw

TheSmartOne (thesmartone):

np